OpenStudy (ny,ny):

The measures of the three angles of a triangle are in the ratio of 2:4:6. What is the measure of the smallest angle of the triangle?

ganeshie8 (ganeshie8):

\(\large \frac{2}{2+4+6} (180)\)

OpenStudy (ny,ny):

How'd you get that?

ganeshie8 (ganeshie8):

angle sum property gives u : sum of angles = 180

ganeshie8 (ganeshie8):

since the angles are in ratio 2 : 4 : 6, the smallest angle wud be 2/(2+4+6) th of the entire sum

ganeshie8 (ganeshie8):

if above doesnt make sense, try another method

ganeshie8 (ganeshie8):

say, the angles are : 2x, 4x and 6x

ganeshie8 (ganeshie8):

by angle sum property they must add up to 180 : 2x + 4x + 6x = 180 solve x

OpenStudy (ny,ny):

x = 15..

OpenStudy (ny,ny):

what does x represent?

ganeshie8 (ganeshie8):

x = 15, so the smallest angle is 2x = 2*15 = 30

ganeshie8 (ganeshie8):

x is just the common factor 30 : 60 : 90

ganeshie8 (ganeshie8):

if u cancel out 15's, u wud get : 2 : 4 : 6

ganeshie8 (ganeshie8):

if u further cancel out 2's, u wud get : 1 : 2 :3

ganeshie8 (ganeshie8):

in ratios, 1 : 2 :3 is same as 2 : 4 : 6 is same as 3 : 12 : 18 ...

ganeshie8 (ganeshie8):

you can think of the previous "x" as a scaling factor

OpenStudy (ny,ny):

okayy, thanks again.

ganeshie8 (ganeshie8):

u wlc!

OpenStudy (ny,ny):

wait so can you do x + 2x + 3x = 180?

OpenStudy (ny,ny):

as the same thing as 2:4:6

ganeshie8 (ganeshie8):

corrected a typo : in ratios, 1 : 2 :3 is same as 2 : 4 : 6 is same as \(\color{red}{6}\): 12 : 18 ...

ganeshie8 (ganeshie8):

of course yes !

ganeshie8 (ganeshie8):

that gives u the same answer

ganeshie8 (ganeshie8):

x + 2x + 3x = 180 solve x

OpenStudy (ny,ny):

30

ganeshie8 (ganeshie8):

so, the smallest angle = x = 30

OpenStudy (ny,ny):

ohhhhh

OpenStudy (ny,ny):

okay that makes sense.

OpenStudy (ny,ny):

youre awesome :)

ganeshie8 (ganeshie8):

haha no you're awesome !

OpenStudy (ny,ny):

lol well thank you.